# Low-Altitude Wireless Networks: A Comprehensive Survey

### v2

@@ 3.1 Timeliness and Latency
In practice, LAWNs are frequently deployed to support time-sensitive applications, where the transmission delay may lead to severe operational and safety implications. To name a few, in emergency response, the timely delivery of situational data such as survivor locations, fire spread, or hazardous material detection is essential for enabling rapid decision-making and coordinated rescue actions. Therefore, one of the core design principles for LAWNs is the minimization of end-to-end latency while maintaining reliable data transmission. To achieve this, finite-block length (FBL) transmission has emerged as an important technique for managing communication in constrained networks. Unlike the traditional Shannon capacity framework, which assumes infinitely long codewords, FBL transmission provides a more practical characterization of the achievable rates given block lengths and error probability. More importantly, the channel conditions in LAWNs often change rapidly, leading to the reliable infinite block length transmission assumption being invalid. The authors of [zhou2023finite, zhou2017second, zhou2019non] investigated finite blocklength multiterminal lossy source coding, providing new benchmarks for efficient compression with multiple encoders, decoders, and source sequences, thereby strengthening the theoretical foundation of multiterminal low-latency communications and greatly promoting the development of low-altitude low-latency networking. The work of [9126811] investigated IoT service provisioning by leveraging FBL transmission in reducing communication delays within drone-aided networks, with drones being both communication relays and computational resources. Moreover, Jin et al. proposed the co-design of sensing, communication, and control in LAWNs, where FBL transmission is utilized with remote control tasks to minimize latency. The results show that optimizing resource allocation for communication and control can lead to lower delays and better overall performance [11045436].
Beyond FBL transmission, some recent studies focus on low-latency communications. For instance, the authors in [liu2023predictive] presented a predictive precoder design for ultra-reliable low-latency communication (URLLC) systems, where a deep learning (DL)-based framework was developed to predict the precoder via historical channel data. By leveraging a convolutional long short-term memory (CLSTM) network, reliable communication with minimal latency can be guaranteed, which demonstrates superior frame error rate (FER) performance approaching the theoretical lower bound with reduced latency. More relatively, the work of [cai2022resource] explored resource allocation for a two-way UAV relaying system under URLLC constraints, focusing on minimizing latency in both the forward and backward links. An iterative approach was proposed to maximize the system throughput while fulfilling the latency and decoding error probability requirements.
@@ 3.2 Energy-Efficient Design
The sustainability of LAWNs is inherently constrained by the limited onboard energy and computational resources of UAVs, making energy-efficient design essential for long-term viability in real-world applications. Early studies primarily examined energy-aware deployment strategies to extend system lifetime. For instance, [9126212] proposed energy-saving deployment algorithms for UAV swarms to optimize wireless coverage while minimizing energy consumption. In a similar approach, the study of [8364586] addressed the rapid deployment problem of heterogeneous UAVs to provide wireless coverage over a target area, where the objective for minimizing the maximum deployment delay to ensure fairness and minimizing the total deployment delay to enhance efficiency were addressed, respectively. However, these works largely focus on stationary optimization and do not fully exploit the high-mobility characteristics of LAWNs. In the context, the authors in [8068199] investigated joint trajectory and transmit power control for UAV relay networks to minimize the outage probability. By carefully devising UAV trajectories and power allocation, energy consumption can be reduced without sacrificing communication reliability. Further, the study in [yang2019energy] proposed a UAV-assisted backscatter communication system for energy-efficient IoT, which shows a tradeoff between collection location and outage probability on the improvement of EE. On this basis, the work of [11129138] investigated the trade-off between communication/sensing performance and energy consumption in ISAC-LAWN systems, where the goal was to maximize worst-case EE by jointly designing beamforming and UAV trajectory.
@@
Note that [9126212, 8364586, 8068199, yang2019energy, 11129138] primarily focused on the transmit power control, significantly lower than the propulsion-related power. Hence, the work of [10159441] proposed a LAWN system simultaneously sensing ground users and forwarding the sensed information. The authors formulated a multi-objective optimization problem to maximize the EE by jointly optimizing user scheduling, movement-related power, and UAV trajectory. While existing methods have improved network EE, their optimization objectives are often limited to reducing the energy consumption of individual UAVs, overlooking the potential gains from cooperative flight dynamics in LAWNs. In real-world deployments, UAVs inevitably interact with surrounding airflow patterns, which can be strategically exploited to boost overall EE. Specifically, airflow over UAV wings generates wake vortices that create an upwash region behind the wingtips. A trailing UAV entering this region experiences reduced aerodynamic drag, thereby lowering propulsion energy consumption, as well-documented in avian flocking. Building on this principle, the studies in [weimerskirch2001energy, wu2025toward] introduced an ISAC-LAWN design that jointly takes into account aerodynamic factors and sensing/communication performance, showing that a ’V’-shaped formation is likely to be aerodynamically favorable.
@@ 3.3 Scalability and Connectivity
Benefiting from the high mobility and on-demand deployment capabilities, scalability and connectivity are essential for LAWNs to adapt to various practical requirements. The performance of large-scale LAWN networks is heavily influenced by the ability to manage communication capacity, coverage area, and resilience under dynamic conditions, such as UAV movement and interference. For instance, to reveal the fundamental performance limits of beam alignment for mobile targets such as UAVs in low-altitude networks, the authors in [zhou2021resolution, zhou2022resolution] investigated mmWave multiple-antenna communications, deriving theoretical benchmarks for optimal non-adaptive query procedures in multidimensional target search and extending them to scenarios with multiple and moving targets. They further demonstrated the benefits of adaptive query procedures, thereby providing finite-blocklength benchmarks that enable reliable connectivity in low-altitude networks. As evidenced by [10557642], wherein network localization and formation control in LAWN systems with asynchronous agents were investigated. Specifically, the authors developed a message-passing algorithm using Gaussian messages for efficient localization and clock offset estimation to maintain robust information exchange across all LAWN nodes. Subsequently, two types of control policies, namely angle-based and distance-based formation control, were employed for adaptive obstacle avoidance. By doing so, the LAWN systems are capable of quickly responding to environmental changes, and thereby, enhancing scalability and connectivity.
@@
However, the coexistence of heterogeneous LAWN nodes sharing spectral and spatial domains creates a complex interference environment that challenges large-scale deployment and reliable interconnection. The main challenges stem from beam pattern overlaps and waveform interactions across both spatial and temporal dimensions. To tackle these issues, the work in [WU2025103451] introduced a mode-switching strategy, where UAVs alternate between jamming and information transmission. By optimizing mode selection, scheduling, and trajectories, this approach ensures strong A2G links even in the presence of mobile ground users. Furthermore, [9143143] proposed a deep learning–based predictive beamforming scheme that mitigated UAV jittering by forecasting the angles between UAV and user, thereby enabling robust and reliable UAV communications. Moreover, the authors in [9551699] proposed a dynamic interference management framework by jointly optimizing UAV trajectories and power allocation, thereby improving overall network efficiency under dense deployments. Beyond intra-LAWN coordination, such solutions can also be extended to satellite integration. For example, the authors in [10680056] designed robust beamforming schemes for SAGIN to guarantee reliable downlink communications.
@@ 4.1 Integrated Sensing and Communications
ISAC represents a transformative paradigm that unifies wireless communications and environmental sensing into a single framework [liu2022integrated, mao1, mao2, wu2025toward1, mao3, mao4]. By adopting a co-design architecture, ISAC eliminates the need for deploying separate sensing or communication subsystems on aerial platforms within LAWNs. This integration not only reduces the spectral footprint, hardware redundancy, and energy consumption, but also enhances the agility of the system. For instance, a UAV transmitting communication signals to ground users can simultaneously leverage the backscattered echoes for obstacle detection or trajectory tracking of neighboring UAVs [jiang2024uav], thereby avoiding reliance on dedicated radar or LiDAR sensors.
@@
This evolutionary trajectory highlights the maturation of ISAC from a dual-functional system into a network-wide intelligent infrastructure, which is particularly crucial for supporting the burgeoning LAE. In such scenarios, UAV networks serve as a vital enabler for applications ranging from urban logistics and aerial inspections to environmental monitoring and emergency response. Through cooperative sensing, these UAV networks can significantly enhance angular resolution and achieve high-fidelity environmental imaging, thereby meeting the stringent requirements of safety, reliability, and responsiveness in low-altitude economic activities. For example, the framework in [wang2025transmit] demonstrated how joint optimization can improve target detection in cluttered environments while mitigating multi-user interference, thereby boosting both sensing accuracy and communication reliability. Moreover, ISAC inherently enables implicit channel state information acquisition through environmental sensing, reducing dependence on conventional pilot-based estimation and further improving spectral efficiency [pyo2025robust, 10791452, 10543024]. This efficiency gain not only conserves limited spectrum resources but also ensures robust connectivity for diverse services in the LAE.
@@
In addition to its dual communication and sensing capabilities, ISAC fundamentally expands the operational capabilities of UAV networks by enabling advanced spatial awareness and autonomy, which is a key technological driver for the LAE. Applications such as high-precision three-dimensional (3D) mapping, dynamic trajectory planning, and swarm coordination become feasible when sensing and communication are tightly integrated. For instance, [lu2024deep] introduced a cooperative uplink-downlink reconstruction framework, where base stations performed active sensing in the downlink while passive sensing was realized through base station-user equipment cooperation in the uplink. The fusion of complementary sensing results yielded higher spatio-temporal resolution, broader coverage, and improved reconstruction accuracy. Similarly, environmental data generated via ISAC was shown to enhance real-time collision avoidance and collaborative perception in UAV swarms [wang2024cooperative], supporting safer autonomous operations in dense and dynamic LAAs. Advanced fusion strategies, such as the multilevel weighted fusion framework in [zhang2025novel], further improved UAV localization by combining symbol-level and data-level information, thereby enhancing estimation performance. Moreover, [hu2025collaborative] proposed a collaborative ISAC-based framework for UAV-terrestrial base station cooperation, achieving high-precision localization alongside reliable communication in scenarios with multiple moving users. Plugging ISAC into LAWNs not only addresses the challenges of spectral congestion and payload constraints. It also enables distributed sensing, collaborative perception, and intelligent UAV autonomy. This progression underscores ISAC’s pivotal role in transforming LAWNs into intelligent, resilient, and spectrum-efficient infrastructures that can support mission-critical low-altitude applications such as urban monitoring, disaster response, traffic management, and environmental surveillance.
@@ 4.2 Networked Control
Networked control serves as a cornerstone of distributed cyber-physical systems, enabling the seamless integration of communication and control functions through shared network infrastructures [meng2024communication, jin2025predictive]. By facilitating real-time exchange of sensory information and control commands among distributed nodes, it provides a scalable mechanism for coordinated operation and decision-making.
@@
This evolutionary trajectory highlights how LAWNs transform rigid and isolated UAV control architectures into intelligent, resilient, and cooperative swarm systems. Such advancements are instrumental in supporting the LAE, where networked UAV swarms serve as the backbone for mission-critical LAE applications including urban surveillance, dynamic aerial coverage, and logistics distribution [mozaffari2019tutorial], all of which require precise synchronization and adaptive coordination in complex low-altitude environments. Through real-time sharing of control signals and sensory information, each UAV dynamically adjusts its trajectory and task allocation based on both local observations and global swarm feedback, thereby ensuring efficient, safe, and scalable operations across diverse LAE scenarios.
This conceptual evolution is further supported by empirical research, where LAWNs-enabled networked control has been shown to improve robustness and adaptability in real-world conditions, providing essential technical support for reliable operations in LAE scenarios. For instance, [chen2025efficient] proposed a multi-UAV cooperative framework specifically designed for dynamic electromagnetic environments, leveraging UAV mobility, adaptive coverage, and reliable line-of-sight connectivity to enhance perception and overall system robustness. Complementing this approach, closed-loop feedback mechanisms were shown to significantly improve swarm adaptability under interference, uncertainty, and partial node failures [qiao2022communication]. Furthermore, [ping2024communication] introduced a distributed optimal control framework for multi-UAV systems that simultaneously supported formation management, obstacle avoidance, and connectivity maintenance while adaptively optimizing communication resources.
Beyond robustness and adaptability, recent studies highlighted the benefits of joint optimization across sensing, communication, and control to achieve end-to-end quality of service (QoS), a capability that can directly enhance mission reliability in critical LAE applications. For instance, [wang2023qos] proposed a joint optimization framework for UAV-based positioning, where UAVs acted as aerial anchors to ensure reliable localization in challenging environments. This approach reduced service failure rates by over $75\%$ and resource consumption by $80\%$, demonstrating the efficiency of cross-domain integration. Air-ground collaboration was also explored in [chen2022formation], where a formation control strategy incorporating spatial synergy enabled UAVs to provide robust relay communication while maintaining coordinated positioning. Likewise, [zhang2020unmanned] presented a lightweight, UAV-based emergency communication framework employing commercial micro-UAVs. By integrating body-conformal omnidirectional antennas with optimized formation strategies, the system enhanced cooperative transmission and ensured reliable coverage in harsh terrains such as mountainous canyons.

### v3

@@ 3.1 Timeliness and Latency
In practice, LAWNs are frequently deployed to support time-sensitive applications, where the transmission delay may lead to severe operational and safety implications. To name a few, in emergency response, the timely delivery of situational data such as survivor locations, fire spread, or hazardous material detection is essential for enabling rapid decision-making and coordinated rescue actions. Therefore, one of the core design principles for LAWNs is the minimization of end-to-end latency while maintaining reliable data transmission. To achieve this, finite-block length (FBL) transmission has emerged as an important technique for managing communication in constrained networks. Unlike the traditional Shannon capacity framework, which assumes infinitely long codewords, FBL transmission provides a more practical characterization of the achievable rates given block lengths and error probability. More importantly, the channel conditions in LAWNs often change rapidly, leading to the reliable infinite block length transmission assumption being invalid. The authors of [207, 209, 208] investigated finite blocklength multiterminal lossy source coding, providing new benchmarks for efficient compression with multiple encoders, decoders, and source sequences, thereby strengthening the theoretical foundation of multiterminal low-latency communications and greatly promoting the development of low-altitude low-latency networking. The work of [193] investigated IoT service provisioning by leveraging FBL transmission in reducing communication delays within drone-aided networks, with drones being both communication relays and computational resources. Moreover, Jin et al. proposed the co-design of sensing, communication, and control in LAWNs, where FBL transmission is utilized with remote control tasks to minimize latency. The results show that optimizing resource allocation for communication and control can lead to lower delays and better overall performance [74].
Beyond FBL transmission, some recent studies focus on low-latency communications. For instance, the authors in [103] presented a predictive precoder design for ultra-reliable low-latency communication (URLLC) systems, where a deep learning (DL)-based framework was developed to predict the precoder via historical channel data. By leveraging a convolutional long short-term memory (CLSTM) network, reliable communication with minimal latency can be guaranteed, which demonstrates superior frame error rate (FER) performance approaching the theoretical lower bound with reduced latency. More relatively, the work of [27] explored resource allocation for a two-way UAV relaying system under URLLC constraints, focusing on minimizing latency in both the forward and backward links. An iterative approach was proposed to maximize the system throughput while fulfilling the latency and decoding error probability requirements.
@@ 3.2 Energy-Efficient Design
The sustainability of LAWNs is inherently constrained by the limited onboard energy and computational resources of UAVs, making energy-efficient design essential for long-term viability in real-world applications. Early studies primarily examined energy-aware deployment strategies to extend system lifetime. For instance, [198] proposed energy-saving deployment algorithms for UAV swarms to optimize wireless coverage while minimizing energy consumption. In a similar approach, the study of [197] addressed the rapid deployment problem of heterogeneous UAVs to provide wireless coverage over a target area, where the objective for minimizing the maximum deployment delay to ensure fairness and minimizing the total deployment delay to enhance efficiency were addressed, respectively. However, these works largely focus on stationary optimization and do not fully exploit the high-mobility characteristics of LAWNs. In the context, the authors in [196] investigated joint trajectory and transmit power control for UAV relay networks to minimize the outage probability. By carefully devising UAV trajectories and power allocation, energy consumption can be reduced without sacrificing communication reliability. Further, the study in [182] proposed a UAV-assisted backscatter communication system for energy-efficient IoT, which shows a tradeoff between collection location and outage probability on the improvement of EE. On this basis, the work of [59] investigated the trade-off between communication/sensing performance and energy consumption in ISAC-LAWN systems, where the goal was to maximize worst-case EE by jointly designing beamforming and UAV trajectory.
@@
Note that [198, 197, 196, 182, 59] primarily focused on the transmit power control, significantly lower than the propulsion-related power. Hence, the work of [106] proposed a LAWN system simultaneously sensing ground users and forwarding the sensed information. The authors formulated a multi-objective optimization problem to maximize the EE by jointly optimizing user scheduling, movement-related power, and UAV trajectory. While existing methods have improved network EE, their optimization objectives are often limited to reducing the energy consumption of individual UAVs, overlooking the potential gains from cooperative flight dynamics in LAWNs. In real-world deployments, UAVs inevitably interact with surrounding airflow patterns, which can be strategically exploited to boost overall EE. Specifically, airflow over UAV wings generates wake vortices that create an upwash region behind the wingtips. A trailing UAV entering this region experiences reduced aerodynamic drag, thereby lowering propulsion energy consumption, as well-documented in avian flocking. Building on this principle, the studies in [164, 166] introduced an ISAC-LAWN design that jointly takes into account aerodynamic factors and sensing/communication performance, showing that a ’V’-shaped formation is likely to be aerodynamically favorable.
@@ 3.3 Scalability and Connectivity
Benefiting from the high mobility and on-demand deployment capabilities, scalability and connectivity are essential for LAWNs to adapt to various practical requirements. The performance of large-scale LAWN networks is heavily influenced by the ability to manage communication capacity, coverage area, and resilience under dynamic conditions, such as UAV movement and interference. For instance, to reveal the fundamental performance limits of beam alignment for mobile targets such as UAVs in low-altitude networks, the authors in [206, 205] investigated mmWave multiple-antenna communications, deriving theoretical benchmarks for optimal non-adaptive query procedures in multidimensional target search and extending them to scenarios with multiple and moving targets. They further demonstrated the benefits of adaptive query procedures, thereby providing finite-blocklength benchmarks that enable reliable connectivity in low-altitude networks. As evidenced by [188], wherein network localization and formation control in LAWN systems with asynchronous agents were investigated. Specifically, the authors developed a message-passing algorithm using Gaussian messages for efficient localization and clock offset estimation to maintain robust information exchange across all LAWN nodes. Subsequently, two types of control policies, namely angle-based and distance-based formation control, were employed for adaptive obstacle avoidance. By doing so, the LAWN systems are capable of quickly responding to environmental changes, and thereby, enhancing scalability and connectivity.
@@
However, the coexistence of heterogeneous LAWN nodes sharing spectral and spatial domains creates a complex interference environment that challenges large-scale deployment and reliable interconnection. The main challenges stem from beam pattern overlaps and waveform interactions across both spatial and temporal dimensions. To tackle these issues, the work in [168] introduced a mode-switching strategy, where UAVs alternate between jamming and information transmission. By optimizing mode selection, scheduling, and trajectories, this approach ensures strong A2G links even in the presence of mobile ground users. Furthermore, [187] proposed a deep learning–based predictive beamforming scheme that mitigated UAV jittering by forecasting the angles between UAV and user, thereby enabling robust and reliable UAV communications. Moreover, the authors in [133] proposed a dynamic interference management framework by jointly optimizing UAV trajectories and power allocation, thereby improving overall network efficiency under dense deployments. Beyond intra-LAWN coordination, such solutions can also be extended to satellite integration. For example, the authors in [113] designed robust beamforming schemes for SAGIN to guarantee reliable downlink communications.
@@ 4.1 Integrated Sensing and Communications
ISAC represents a transformative paradigm that unifies wireless communications and environmental sensing into a single framework [104, 93, 189, 170, 190, 112]. By adopting a co-design architecture, ISAC eliminates the need for deploying separate sensing or communication subsystems on aerial platforms within LAWNs. This integration not only reduces the spectral footprint, hardware redundancy, and energy consumption, but also enhances the agility of the system. For instance, a UAV transmitting communication signals to ground users can simultaneously leverage the backscattered echoes for obstacle detection or trajectory tracking of neighboring UAVs [73], thereby avoiding reliance on dedicated radar or LiDAR sensors.
@@
This evolutionary trajectory highlights the maturation of ISAC from a dual-functional system into a network-wide intelligent infrastructure, which is particularly crucial for supporting the burgeoning LAE. In such scenarios, UAV networks serve as a vital enabler for applications ranging from urban logistics and aerial inspections to environmental monitoring and emergency response. Through cooperative sensing, these UAV networks can significantly enhance angular resolution and achieve high-fidelity environmental imaging, thereby meeting the stringent requirements of safety, reliability, and responsiveness in low-altitude economic activities. For example, the framework in [152] demonstrated how joint optimization can improve target detection in cluttered environments while mitigating multi-user interference, thereby boosting both sensing accuracy and communication reliability. Moreover, ISAC inherently enables implicit channel state information acquisition through environmental sensing, reducing dependence on conventional pilot-based estimation and further improving spectral efficiency [131, 169, 199]. This efficiency gain not only conserves limited spectrum resources but also ensures robust connectivity for diverse services in the LAE.
@@
In addition to its dual communication and sensing capabilities, ISAC fundamentally expands the operational capabilities of UAV networks by enabling advanced spatial awareness and autonomy, which is a key technological driver for the LAE. Applications such as high-precision three-dimensional (3D) mapping, dynamic trajectory planning, and swarm coordination become feasible when sensing and communication are tightly integrated. For instance, [107] introduced a cooperative uplink-downlink reconstruction framework, where base stations performed active sensing in the downlink while passive sensing was realized through base station-user equipment cooperation in the uplink. The fusion of complementary sensing results yielded higher spatio-temporal resolution, broader coverage, and improved reconstruction accuracy. Similarly, environmental data generated via ISAC was shown to enhance real-time collision avoidance and collaborative perception in UAV swarms [151], supporting safer autonomous operations in dense and dynamic LAAs. Advanced fusion strategies, such as the multilevel weighted fusion framework in [191], further improved UAV localization by combining symbol-level and data-level information, thereby enhancing estimation performance. Moreover, [65] proposed a collaborative ISAC-based framework for UAV-terrestrial base station cooperation, achieving high-precision localization alongside reliable communication in scenarios with multiple moving users. Plugging ISAC into LAWNs not only addresses the challenges of spectral congestion and payload constraints. It also enables distributed sensing, collaborative perception, and intelligent UAV autonomy. This progression underscores ISAC’s pivotal role in transforming LAWNs into intelligent, resilient, and spectrum-efficient infrastructures that can support mission-critical low-altitude applications such as urban monitoring, disaster response, traffic management, and environmental surveillance.
@@ 4.2 Networked Control
Networked control serves as a cornerstone of distributed cyber-physical systems, enabling the seamless integration of communication and control functions through shared network infrastructures [118, 75]. By facilitating real-time exchange of sensory information and control commands among distributed nodes, it provides a scalable mechanism for coordinated operation and decision-making.
@@
This evolutionary trajectory highlights how LAWNs transform rigid and isolated UAV control architectures into intelligent, resilient, and cooperative swarm systems. Such advancements are instrumental in supporting the LAE, where networked UAV swarms serve as the backbone for mission-critical LAE applications including urban surveillance, dynamic aerial coverage, and logistics distribution [120], all of which require precise synchronization and adaptive coordination in complex low-altitude environments. Through real-time sharing of control signals and sensory information, each UAV dynamically adjusts its trajectory and task allocation based on both local observations and global swarm feedback, thereby ensuring efficient, safe, and scalable operations across diverse LAE scenarios.
This conceptual evolution is further supported by empirical research, where LAWNs-enabled networked control has been shown to improve robustness and adaptability in real-world conditions, providing essential technical support for reliable operations in LAE scenarios. For instance, [31] proposed a multi-UAV cooperative framework specifically designed for dynamic electromagnetic environments, leveraging UAV mobility, adaptive coverage, and reliable line-of-sight connectivity to enhance perception and overall system robustness. Complementing this approach, closed-loop feedback mechanisms were shown to significantly improve swarm adaptability under interference, uncertainty, and partial node failures [132]. Furthermore, [128] introduced a distributed optimal control framework for multi-UAV systems that simultaneously supported formation management, obstacle avoidance, and connectivity maintenance while adaptively optimizing communication resources.
Beyond robustness and adaptability, recent studies highlighted the benefits of joint optimization across sensing, communication, and control to achieve end-to-end quality of service (QoS), a capability that can directly enhance mission reliability in critical LAE applications. For instance, [162] proposed a joint optimization framework for UAV-based positioning, where UAVs acted as aerial anchors to ensure reliable localization in challenging environments. This approach reduced service failure rates by over $75\%$ and resource consumption by $80\%$, demonstrating the efficiency of cross-domain integration. Air-ground collaboration was also explored in [32], where a formation control strategy incorporating spatial synergy enabled UAVs to provide robust relay communication while maintaining coordinated positioning. Likewise, [195] presented a lightweight, UAV-based emergency communication framework employing commercial micro-UAVs. By integrating body-conformal omnidirectional antennas with optimized formation strategies, the system enhanced cooperative transmission and ensured reliable coverage in harsh terrains such as mountainous canyons.